OpenStudy (anonymous):

circumference when diameter is 8.4cm

OpenStudy (johnweldon1993):

You know that diameter...is 2 times the radius C = 2piR So half your diameter to find your Radius....and plug that in for 'R' and use that equation to solve for Circumference!

OpenStudy (raden):

or C = pi * d (with d is the diameter)

OpenStudy (anonymous):

I think for beginner , it is helpful to just remember one formula

OpenStudy (texaschic101):

C = (pi)(d) C = (3.14)(8.4) can you take it from here ?